OpenStudy (anonymous):

Le Man had $25. After buying 3 adult tickets and 4 children tickets to a concert, she had some money left. If she used the remaining money to buy a child's ticket, she would have 50 cents left. If she used the remaining money to buy an adults children instead of a child's ticket she would be short of $1. How much did an adult ticket and a child's ticket cost?

OpenStudy (anonymous):

Let the price of adult tickets be a Let the price of children tickets be c 25 - (3a + 4c) - c = 0.5 25 - 3a - 5c = 0.5 24.5 - 5c = 3a a = (24.5 - 5c)/3 25 - (3a + 4c) - a = -1 25 - 4a - 4c = -1 26 = 4a + 4c 26 = 4 (24.5 - 5c)/3 + 4c Solve for c

OpenStudy (anonymous):

c = 2.5 a = 4

OpenStudy (anonymous):

that works. you could also put \[25=3a+5c+.5\iff 24.5=3a+5c\] \[25=4a+4c-1\iff 26=4a+4c\] and solve the system